THE WEST COAST OPTIMIZATION MEETING

Depts. of Mathematics and Applied Mathematics, University of Washington All talks will be in Guggenheim 317

FRIDAY, NOVEMBER 1

6:30–9:30 Party at Terry Rockafellar’s home, 4531 NE 93rd Street, 206–527–9637 The cost per person will be $10/“student” and $15/“others.”

SATURDAY, NOVEMBER 2

8:30–9:00 —————Refreshments in Guggenheim 408, the Applied Math Lounge

9:00–9:35 “Merit Functions for Semi-definite Complementarity Problems,” Paul

Tseng, University of Washington, Seattle 9:35–10:00 “Dynamic Splitting: an Operator Splitting Algorithm for the Multistage

Stochastic Programming Problem,” Dave Salinger, Ph.D. candidate (U. of W.)

10:00–10:20 ————Refreshments in Guggenheim 408, the Applied Math Lounge

10:20–10:55 “On the Certainty Equivalence Principle in Nonlinear Control, and the Hamilton-Jacobi-Bellman-Isaac Equation,” Andrei Vityaev, University of Washington, Seattle

10:55–11:20 “Interior and Non-Interior Path-Following Methods for Complementarity Problems,” Song Xu, Ph.D. candidate (U. of W.)

11:20–11:30 ————Refreshments in Guggenheim 408, the Applied Math Lounge

11:30–12:05 “A Stochastic Branch-and-Bound Method,” Andrzej Ruszczynski, University of Wisconsin, Madison

12:05–13:30 ————Lunch expedition to University Avenue

13:30–14:05 “Necessary Conditions for NonLipschitzian Optimal Control Prob-

lems,” Philip Loewen, University of British Columbia, Vancouver 14:05–14:40 “Proximal Analysis and the Minimal Time Function,” Peter Wolenski,

Louisiana State University, Baton Rouge

14:50–15:00 ————Refreshments in Guggenheim 408, the Applied Math Lounge

15:00–15:35 “Approximate Jacobians and Nonsmooth Analysis,” V. Jeyakumar, University of New South Wales, Sidney

The West Coast Optimization Meeting occurs twice each year. Contact: Prof. J. M. Borwein at the Dept. of Mathematics and Statistics, Simon Fraser

University, Vancouver: (604) 291-3070, e-mail jborwein@cecm.sfu.ca Prof. R. T. Rockafellar at the Dept. of Mathematics, University of Washington,

Seattle: (206) 543-1916, e-mail rtr@math.washington.edu